# CMAT: A Multi-Agent Collaboration Tuning Framework for Enhancing Small Language Models

### v4

@@ 4.4 Results analysis
The results in Table 2 underscore the effectiveness of our fine-tuning methods, especially for the TinyAgent models. Tinyagent-1.8B demonstrates significant performance in the KG task, on par with advanced models like GPT-3.5. Tinyagent-7B also showcases its strengths, notably in the DB task, where it surpasses its foundational model Antonello et al. (2020), CodeLlama-7B, and offers competitive scores against GPT-4. These findings indicate the TinyAgent models’ capacity to match or even surpass models with larger parameters in certain aspects. Moreover, the CMAT framework’s potential to enhance the capabilities of smaller-scale models is highlighted, allowing the TinyAgent models to closely compete with the performance of advanced models such as GPT-4.
@@
As illustrated in Figure 1, Our comparative analysis indicates that Tinyagent models, refined from Qwen-1.8B and CodeLlama-7B, exhibit superior performance to their base models. The incorporation of the CMAT framework further amplifies their functionality, equipping these small Models to match the capabilities of GPT-3.5. This performance boost is credited to CMAT’s optimization of model interactions and its strategic use of memory modes for specific tasks, confirming its effectiveness in enhancing the sophistication of fine-tuned models Deshpande et al. (2021).
4.5 Error analysis
@@
4.6 Ablation Study
The Table 3 presents an ablation study on the TinyAgent-7B model, delineating the impact of agent-specific and general instructions on task performance. The composite model, TinyAgent-7B, demonstrates the highest efficacy, notably in WS and DB tasks, which implies its adeptness in handling complex e-commerce interactions and database management. The agent-only variant exhibits a decline in performance, suggesting that while task-specific instructions are crucial, they are not wholly sufficient for the breadth of tasks such as KG. The general-only model’s performance is considerably reduced across all tasks, with a complete inability to perform in KG and ALF, highlighting the indispensability of agent-specific instructions. This data underscores the necessity of integrating both agent-specific and general instructions to enhance the versatility and effectiveness of AI models in diverse task domains.
5 Conclusions
Limitations
In this study, we demonstrated the potential for performance improvement by applying the CMAT framework to TinyAgent series models and other large language models (LLMs). However, there are clear limitations to the research: First, although most models showed improved performance, some models saw limited improvement due to weaker base agent capabilities, indicating that the effectiveness of the CMAT framework might vary significantly between different models; second, the limitations of datasets and task types could affect the broad applicability of the conclusions, while low-quality datasets could negatively impact model performance; lastly, although evaluations based on AgentBench ensured fairness, they might not fully reflect the complexity of real-world scenarios, and due to computational resource constraints, larger-scale models could not be tested. This underscores the importance of future work to consider a wider range of models, datasets, and task types, especially the implementation of optimization strategies and framework applications in resource-constrained situations.
Appendix A Implementation Settings
In this paper, we describe an experiment conducted using the Low-Rank Adaptation (LoRA) fine-tuning method to enhance the performance of various models Wang et al. (2005). The accuracy of the LoRA method is of paramount importance in dealing with personalized and emotionally rich content. It enables the models to adapt to new data features while maintaining their core capabilities Bai et al. (2015).
During the experiment, we set the temperature parameter of the test models to 0.7 to increase the diversity of the content, and adjusted the top-p value to 0.95 to improve the precision of the generated content. We employed a learning rate of 2e-4 and beta values of (0.9, 0.999) to ensure the stability of the training process. The batch size was set to 4, with gradient accumulation, to ensure efficiency within the limits of computational resources. To balance innovation and coherence, we used LoRA parameters with a rank of 8 and an alpha value of 32, and adjusted both the top-p value and the temperature parameter to 0.7.These adjustments significantly enhanced the models’ flexibility and accuracy in handling personalized and emotionally rich content.
| OS | DB | KG | ALF | WS | W2M | |
@@ Appendix A Implementation Settings
Appendix B Evaluation criteria
Operating systems Integrating LLMs into operating systems offers vast potential for automating and optimizing tasks. This integration demands a secure, user-friendly interface for effective LLM-OS interaction and requires LLMs to accurately understand the OS context for informed operations. Ensuring the safety of these operations is paramount to prevent misuse. Moreover, the system must handle errors and provide clear feedback to users, enhancing interaction and control. Addressing these aspects can revolutionize computer interaction and efficiency across industries.
Database Database (DB). Due to the crucial and challenging nature of database analysis in many daily affairs, it is paramount to examine the abilities of LLMs to operate on real databases via SQL. Previous research has placed significant emphasis on individual procedures, such as showcasing the effectiveness of LLMs in automating database access through T5QL, a new SQL generation method. Additionally, utilizing fine-tuned LLMs (such as GPT-3.5) to extract and link complex scientific information from scientific texts has demonstrated the capacity of LLMs to obtain structured knowledge from unstructured text and subsequently construct large databases Dunn et al. (2022).
WebShop represents an innovative simulation of an e-commerce website environment, featuring 1.18 million real-world products and 12,087 crowd-sourced text instructions. This platform challenges agents to navigate through multiple types of webpages and perform a variety of actions to find, customize, and purchase products according to given instructions. WebShop presents several challenges, including understanding compositional instructions, query (re-)formulation, dealing with noisy text in webpages, and conducting strategic exploration.
Knowledge Graphs The utilization of LLMs in constructing and interacting with knowledge graphs (KG) offers a promising avenue for enhancing semantic understanding and information retrieval. This involves assessing the models’ ability to not only generate but also interpret complex interrelations within data, facilitating more intuitive and context-aware responses. The effectiveness of LLMs in this domain could significantly improve AI’s capacity for reasoning and decision-making based on structured knowledge.
Mind2Web Mind2Web (M2W) is a dataset for developing web agents that perform complex tasks on real websites via language instructions. It features over 2,000 tasks across 137 sites from 31 domains. M2W’s real web environments and diverse user interactions make it a crucial platform for advancing AI navigation capabilities.
ALFWorld bridges interactive TextWorld environments with embodied tasks from the ALFRED dataset, enabling agents to learn abstract strategies and apply them to real-world tasks. It facilitates abstract reasoning and concrete execution, allowing agents to plan actions in a text-based simulator and then execute these tasks in a visual environment. This approach enhances agent generalization and problem-solving skills across various domains, such as language understanding and visual navigation, by leveraging a modular design that simplifies research improvements.
Appendix C Details about Experiment
| Evaluation Method | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L |
| prompt - High-quality | 44.4 | 57.3 | 35.0 | 42.5 |
| prompt - Low-quality | 15.2 | 27.4 | 10.3 | 16.8 |
| without prompts | 26.8 | 47.2 | 30.2 | 26.7 |
In our research, we focused on exploring the impact of meticulously crafted prompts on the performance of Natural Language Processing (NLP) models in code correction tasks as shown in Table 5. By examining prompts of varying quality—high-quality, low-quality, and no prompts—we aimed to reveal how these factors influence the model’s capability, especially in identifying and correcting programming errors. Our findings indicate that fine-tuned high-quality prompts significantly enhance the model’s ability to write and rigorously check code, enabling it to produce high-quality code and conduct stricter code inspections. As demonstrated in the Table 23 25 .In contrast, low-quality prompts might lead to the model generating nonsensical and ineffective code, and in situations where correct code is provided, the model might even alter it to incorrect code. The decline in model output quality in these instances is primarily due to the vague and unclear descriptions provided by the prompts and the lack of specific guidance, which prevents the model from effectively utilizing the knowledge it acquired during training to focus on key error points and make accurate correction decisions. Moreover, our study also showed that in the absence of any prompt guidance, the model’s performance tends to be limited by the quality and scope of its training data, particularly when faced with novel or complex error types. This finding underscores the importance of high-quality prompts in designing effective NLP models for complex tasks like code correction. High-quality prompts not only guide the model to more accurately understand and address specific programming issues but also reduce noise in the processing, thereby improving overall efficiency and accuracy.
Appendix D Prompts for Tasks
Appendix E Examples of Tasks
All tasks in the dataset are divided into six categories. For each category, we present an example of successful execution as shown in Table 13 22.
DATABASE(DB) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
WEBSHOP(WS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
ALFWORLD(ALF) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
MIND2WEB(M2W) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
OPERATING SYSTEM(OS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
OPERATING SYSTEM(OS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
ALFWORLD(ALF) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
MIND2WEB(M2W) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
MIND2WEB(M2W) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
WEBSHOP(WS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
WEBSHOP(WS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
DATABASE(DB) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
DATABASE(DB) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]

### v6

@@ IV-D Results analysis
The results in Table II underscore the effectiveness of our fine-tuning methods, especially for the TinyAgent models. Tinyagent-1.8B demonstrates significant performance in the KG task, on par with advanced models like GPT-3.5. Tinyagent-7B also showcases its strengths, notably in the DB task, where it surpasses its foundational model [44], CodeLlama-7B, and offers competitive scores against GPT-4. These findings indicate the TinyAgent models’ capacity to match or even surpass models with larger parameters in certain aspects. Moreover, the CMAT framework’s potential to enhance the capabilities of smaller-scale models is highlighted, allowing the TinyAgent models to closely compete with the performance of advanced models such as GPT-4.
@@
As illustrated in Figure 1, Our comparative analysis indicates that Tinyagent models, refined from Qwen-1.8B and CodeLlama-7B, exhibit superior performance to their base models. The incorporation of the CMAT framework further amplifies their functionality, equipping these small Models to match the capabilities of GPT-3.5. This performance boost is credited to CMAT’s optimization of model interactions and its strategic use of memory modes for specific tasks, confirming its effectiveness in enhancing the sophistication of fine-tuned models [45].
Table III presents the impact of different prompting strategies on performance metrics. High-quality prompts significantly outperform low-quality prompts and scenarios without prompts across all evaluation metrics, demonstrating the importance of prompt design in optimizing model performance.
| Evaluation Method | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L |
| prompt - High-quality | 44.4 | 57.3 | 35.0 | 42.5 |
| prompt - Low-quality | 15.2 | 27.4 | 10.3 | 16.8 |
| without prompts | 26.8 | 47.2 | 30.2 | 26.7 |
IV-E Error analysis
@@
As illustrated in Table IV, the distribution of various execution results across six tasks highlights the prevalence of specific error types, such as exceeding task limits (TLE) and invalid actions, which point to limitations in LLM agents’ reasoning and decision-making within constrained timeframes.
| Execution Results | OS | DB | KG | ALF | WS | M2W |
@@
IV-F Ablation Study
The Table V presents an ablation study on the TinyAgent-7B model, delineating the impact of agent-specific and general instructions on task performance. The composite model, TinyAgent-7B, demonstrates the highest efficacy, notably in WS and DB tasks, which implies its adeptness in handling complex e-commerce interactions and database management. The agent-only variant exhibits a decline in performance, suggesting that while task-specific instructions are crucial, they are not wholly sufficient for the breadth of tasks such as KG. The general-only model’s performance is considerably reduced across all tasks, with a complete inability to perform in KG and ALF, highlighting the indispensability of agent-specific instructions. This data underscores the necessity of integrating both agent-specific and general instructions to enhance the versatility and effectiveness of AI models in diverse task domains.
| Models | OS | DB | KG | ALF | WS | M2W |
| tinyagent-7b | 27.3 | 43.0 | 38.0 | 10.0 | 61.8 | 14.0 |
| - agent only | 20.1 | 39.3 | 25.0 | 2.0 | 55.7 | 7.0 |
| - general only | 9.7 | 5.4 | 0.0 | 0.0 | 26.6 | 5.0 |
V Conclusions